# REPS: Recycled Entropy Packet Spraying for Adaptive Load Balancing and Failure Mitigation

### v3

@@ 4.3.3. Network Failures
We focus our attention to cases where the network encounters a failure during operation. We collect data from several previous works on networking failures and also study internal logs to simulate the most commonly reported cases (Gunawi et al., 2018; Singh et al., 2020, 2021a, 2021b). Since the probability of a failure happening during a short simulation is low, we simulate worst case scenarios where we force individual failures to happen. Most particularly we focus on total or partial failures of cables and switches. We note that we limit our failures to components that would not prevent the workload from completing (single point of failure).
@@
We study a simple synthetic case where we fail, at different times, two uplinks of a TOR switch: a shorter $100\;\mu\text{s}$ failure and a longer $300\;\mu\text{s}$ failure. As suggested in Section 3.2, we advise to set the timeout for freezing mode of REPS to a number which is consistent with the average failure duration (which we artificially lower only for this experiment).
In Figure 7, OPS keeps choosing all paths equally (although at lower rate due to CC activation), while REPS once it enters freezing mode, stops selecting the failing path all together after only one timeout period (order of tens of microseconds). Afterwards, once the failure stops, REPS also exits freezing mode and converges once again quickly to use all paths. The overall result is that, compared to OPS, REPS completes the workload more than 35% faster even with such a short failure and, more importantly, reduces the number of dropped packets by 2.5×2.5$\times$2.5 ×. We showcase a similar analysis but for incremental failures in Appendix D.3.
@@
We showcase three cases with a series of failure modes in Figure 6. We start each failure mode after a fixed amount of time to ensure initial CC convergence. We can see when dealing with total failures that REPS provides a dramatic speedup over OPS but also other load balancer algorithms. This is because of freezing mode that helps REPS to quickly converge to a safe configuration after detecting a failure, considerably faster than the time needed for ECMP routing to update to exclude the failing path. Positively, we note that the gains with REPS inrease with the amount of failures. Furthermore, random drops (e.g., because of BER) does not affect negatively REPS performance.
To further demonstrate the resilience of REPS, we evaluate its performance under an extreme scenario characterized by increasingly large and long-lasting network failures during a permutation. As shown in Figure 8, REPS performs close to an ideal load balancer, even with 50% of network cables failing, while PLB, the second-best alternative, significantly lags behind.
@@
4.4. REPS-FPGA Evaluation
@@ 4.4.1. Healthy Symmetric Network Conditions
The first set of results focuses on a baseline healthy symmetric network configuration. Figure 9(a) shows per-flow goodput defined as end-to-end useful bit rate observed by application, after header, overheads, retransmissions, etc. We use OPS and REPS across two experimental configurations: denoted as setup-1 and setup-2. In setup-1, all FPGA endpoints under the two T0s are active while in setup-2, 40 out of 64 FPGA endpoints are active.
We present results from both of these configurations as we observe small unexpected performance variations depending on which and how many switch ports are active in experiments near the peak network performance levels using all switch ports (setup-1). These variations appear to be related to internal switch microarchitectural details such as port-buffer affinity and vendor-specific scheduling policies. The setup-2 uses a subset of the switch ports and eliminates most of these vendor-specific and implementation-related behaviors. To get a better understanding of this behavior we performed a sweep where we capped the TX rate of our FPGA NICs and discovered that the slight degradation for setup-1 when using REPS appears to subside if the TX rate is capped at 95 Gbps.
@@ 4.4.2. Asymmetric Network Conditions
We evaluate the performance of REPS under asymmetric network conditions. We connect 16 endpoints through two T0 switches (8 endpoints each) with a total of 4 links to a pair of T1 switches. To demonstrate the adaptive load balancing capabilities of REPS, we change the link speed of one T0-T1 link from 400 Gbps to 200 Gbps, creating asymmetry in the network. Fig. 9(b) shows the per-flow goodput as observed by the application while Fig. 10(a) the FCT distribution. OPS sends packets across all paths (including those crossing the 200 Gbps link) with equal probability and is ultimately capped by the slower 200 Gbps path. The ECN marking on the 200 Gbps path causes the CC algorithm to throttle all flows and eventually match the capacity of that single slower link, thus leading to underutilization of the remaining 400 Gbps links (that are running at 50% utilization).
REPS can gracefully adapt in such a scenario as the cached entropies will reflect the network asymmetry and result in a path distribution that is skewed to tailor to the relative capacity of the available paths. In this example, REPS can reach high utilization with average per-flow goodput within 5% of the ideal fair-share target.
@@
Figure 9. REPS-FPGA impact on goodput.
@@
Figure 10. REPS-FPGA impact on FCT and packet drops
@@ 4.4.3. Network Failures
We also evaluate the performance of REPS in the presence of network failures. To demonstrate the robustness and resilience of REPS in the context of link failures, Fig. 10(b) shows total packet drops (average across five runs each) observed in a large-scale 128 endpoint run (endpoints split across 2 T0s connected through 8 T1s) where we abruptly bring down a T0-T1 link during the experiment. While the network is trying to recover from the impact of this event (which in our environment can take in the order of 100s of milliseconds), OPS continues sending packets across all paths (including those affected by the link that went down). The freezing capability of REPS can quickly adapt to such events (within the order of an RTO) and avoid sending packets down the affected paths as the entropy cache is replenished from packets traversing remaining healthy paths.
4.5. REPS Applicability
In this section, we briefly evaluate REPS, in simulations, under different scenarios by changing the EVS size, the ACK coalescing ratio and the underlying CC algorithm. Generally, we believe that REPS can work well under many different circumstances, topologies and workloads.
@@ 4.5.1. EVS Size
In Appendix B, we prove that OPS routing, done through EVs and ECMP hashing, requires a large EVS to work correctly and reduce collisions. Moreover, this requirement also grows with the number of output ports in a switch as demonstrated in Section 5.1. On the other hand, REPS, due to its adaptive nature, can drastically reduce the EVS size and still work well. We show this in the left plot of Figure 12 where we compare, in a real scenario, OPS and REPS when using 32, 256, and 64K EVs. REPS works equally well with 256 and 64K EVs and is only 8% slower with 32 EVs. On the other hand, OPS is 21% and 64% slower with 256 and 32 EVs when compared to 64K. This confirms that REPS could potentially work well even just with 1 byte for the EVS. We note that while OPS could be implemented without using EVS—such as through round-robin selection or by making random choices directly at the switch—these approaches introduce additional challenges (Section 5.2).
4.5.2. ACK Coalescing
We have primarily evaluated REPS without ACK coalescing, as this configuration allows REPS to operate with the most up to date data. However, some transport protocols permit ACK coalescing, where the receiver sends an ACK packet only after receiving $n$ data packets from the sender. In theory, the coalesced ACK packet could return all previous non-ECN marked entropies in its header, but we focus on the worst case scenario for REPS where the ACK only carries the entropy of the packet that triggered it.
We run a real 8 MiB permutation workload simulating different ACK coalescing ratios. As we can see in the left part of Figure 11, REPS with 2:1, 4:1 and 8:1 coalescing ratios does significantly better than OPS, while with 16:1 it starts losing its advantage. However, it should be noted that in case of asymmetries or failures (right Figure 11), REPS does much better due to its various improvements over OPS even at 16:1 ratio. We confirm these results theoretically in Appendix D.1.
Finally, we note that the trade-off of sending more ACKs is worth the effort if the underlining hardware supports such rate and the impact on the network traffic is minimal ( 1%) since ACKs are relatively small (64B) compared to the packet size usually used in modern interconnects (4KiB or 8KiB).
@@ 4.5.3. Different CC Algorithms
In principle, REPS has been designed to work with any CC algorithm as long as there is no over-reaction to out-of-order packets and ECN support. In this paper, we have, so far, evaluated REPS working alongside a tuned version of DCTCP. However, as we will see, REPS can work well even with other algorithms. Moreover, we envision also a version of REPS that could work just with delay if ECN is not supported but we do not go into details here. For example, in the right plot of Figure 12, we run a simple 8 MiB permutation workload without failure for DCTCP, EQDS and a proprietary CC algorithm. REPS can help all of these CC algorithms when compared to OPS.
@@ 5. Theoretical Verification
To support our experimental findings, we present a theoretical first principles analysis of OPS and demonstrate how it can lead to arbitrarily long queues. To address this issue, we develop a new recycled balls-into-bins model and prove its local convergence. This model serves as the theoretical intuition behind the REPS protocol (Section 3).
@@ 5.1. Recycled Balls-into-Bins Model
At maximum injection rate, OPS suffers from severe load-imbalance that eventually leads to queues of arbitrary length building up. We explain this behavior with an infinite batched balls-into-bins model (Berenbrink et al., 2018; Becchetti et al., 2019; Los and Sauerwald, 2023). In contrast, we show that recycling good paths such as in REPS leads to a convergent behavior and logarithmically-bounded queues.
@@
For $n\geq 16$, $\tau\geq 4\ln n$, $b\geq 2.4\ln n$, recycled balls-into-bins converges in $O(n\log n)$ expected steps. Every bin has $O(\log n)$ elements throughout, with probability $1-o(1)$.
@@
To visualize these behaviors, we model the balls-into-bins problem with both OPS and with the recycled balls improvement. In Figure 14, we set $n=8$ (for visualization purposes, but this holds true for more realistic $n$ values) and showcase the queues’ evolution for oblivious packet spraying and the recycled balls model. The results confirm our findings that oblivious packet spraying will see the queues grow unbounded while the recycled balls model will eventually converge and keep all queues below the threshold ($\tau$) value. This is also consistent with our previous simulation results (e.g. Figure 1 and Figure 3).
@@ 5.2. Limitations and Alternatives
While the provided models perform well in an idealized setting with maximal injection rates, in a real network scenario, CC algorithms would reduce the sending rate if a large queue were to build up. However, such activations of the CC would result in a slowdown of the sending rate and, hence, an increase of the completion time of a workload as we can see in Section 4.3.1. Moreover, we note that while real scenarios would have more subtle interactions not modelled here (like using distinct $KMin$ and $KMax$), the key results hold and are consistent with our evaluations in Section 4. Finally, we note that here we provide a proof only about the local convergence of REPS and that other models such as static path assignments and round-robin across the ports would also work well and potentially provide no queueing. However, such approaches are not viable in reality for a series of reasons: they cannot react well to partial or total failures, they would struggle with multiple tiers topologies and they require knowing the workload in advance to properly assign paths (Gangidi et al., 2024).
@@ 6. Related Work
The literature for load balancing designs is vast, however most of it focuses on TCP-based deployments where out-of-order packets are not desirable (Benson et al., 2011; Zhang et al., 2017; Al-Fares et al., 2010; Vanini et al., 2017; Dixit et al., 2013). In contrast, REPS aligns with the new transports (UE (Consortium, 2024), Adaptive RoCE by NVIDIA (NVIDIA, 2024), Falcon (goo, 2023)) where out-of-order packets help to leverage the capacity of multi-path networks.
@@
Load balancers with per-packet granularity, such as OPS (Dixit et al., 2013), help drastically reduce ECMP collisions, but still are oblivious to asymmetries. As demonstrated in Section 5 OPS can suffer even in symmetric cases. MPRDMA (Lu et al., 2018) also uses ECN for load balancing like REPS, however it requires probing and ACK clocking and does not offer caching of entropies. Hermes (Zhang et al., 2017) combines both ECN and delay but it works best with TCP-like protocols and has many parameters to tune.
@@
Proteus (Hu et al., 2023) focuses on optimizing the load balancing for lossless PFC networks while REPS focuses on lossy networks.
@@ 7. Conclusion
We presented REPS, a simple, lightweight, yet highly effective load-balancing mechanism designed to meet the constraints of next-generation datacenter networks tailored for AI workloads. As demonstrated in our extensive evaluations, conducted through both simulations and FPGA hardware, REPS’ adaptive entropy caching enhances end-to-end performance across multiple critical metrics, including average flow completion time, runtime, and packet loss. REPS outperforms ECMP and OPS by up to 6x and 1.25x in symmetric networks, and by up to 4.5x and 1.5x in asymmetric networks, outperforming OPS by as much as 100x during short-term transient link failures while reducing packet drops by over 70x. We also showed how REPS can work well under various network configurations demonstrating its flexibility to adapt to different scenarios while remaining light-weight on memory requiring only 25 bytes of state per-connection.
Appendix A Freezing Mode in REPS
Ideally, REPS should enter freezing mode only upon detecting a network failure. To achieve this, we employ two strategies:

### v4

@@ 4.3.3. Network Failures
We focus our attention to cases where the network encounters a failure during operation. We collect data from several previous works on networking failures and also study internal logs to simulate the most commonly reported cases (Gunawi et al., 2018; Singh et al., 2020, 2021). Since the probability of a failure happening during a short simulation is low, we simulate worst case scenarios where we force individual failures to happen during the runtime of our simulations. Most particularly we focus on total or partial failures of cables and switches. We note that we limit our failures to components that would not prevent the workload from completing (single point of failure).
@@
We begin with a synthetic scenario in which two TOR-switch uplink failures occur at different times: one lasting $100\,\mu\mathrm{s}$ starting at $t=100\,\mu\mathrm{s}$, and another lasting $200\,\mu\mathrm{s}$ starting at $t=350\,\mu\mathrm{s}$.
In Figure 6, OPS keeps choosing all paths equally (although at lower rate due to CC activation), while ARCANE, once it enters freezing mode, stops selecting the failing paths all together after only one timeout period (order of tens of microseconds). Afterwards, once the failure stops, ARCANE also exits freezing mode and converges once again quickly to use all paths. The overall result is that, compared to OPS, ARCANE completes the workload more than 35% faster even with such a short failure and, more importantly, reduces the number of dropped packets by 2.5×2.5$\times$2.5 ×. We showcase a similar analysis but for incremental and concurrent failures in Appendix D.3.
@@
We showcase three cases with a series of failure modes in Figure 7. We introduce each failure mode after a fixed interval to highlight how ARCANE dynamically reacts to changing network conditions, and then let the failure persist for the remainder of the experiment. We can see when dealing with total failures that ARCANE provides a dramatic speedup over OPS but also other load balancer algorithms. This is because of freezing mode which helps ARCANE stick to the safe path configuration in its cache, allowing it to react immediately upon failure, much faster than the time ECMP routing takes to converge and exclude the failing path. Positively, we note that the gains with ARCANE increase with the amount of failures. Furthermore, random drops do not affect negatively ARCANE performance.
@@
To further demonstrate the resilience of ARCANE, we evaluate its performance under an extreme and unlikely scenario characterized by increasingly large and long-lasting network failures during a permutation. As shown in Figure 8, ARCANE performs close to an ideal load balancer, even with 50% of network cables failing, while PLB, the second-best alternative, lags significantly behind.
This analysis proves that ARCANE is able to guarantee good load balancing performance under short-lived failures, long-lived failures and a combination of the two even with catastrophically high failure rates.
4.4. ARCANE-FPGA Evaluation
@@ 4.4.1. Healthy Symmetric Network Conditions
The first set of results focuses on a baseline healthy symmetric network configuration. Figure 9(a) shows per-flow goodput defined as end-to-end useful bit rate observed by application, after header, overheads, retransmissions, etc. We use OPS and ARCANE across two experimental configurations: denoted as setup-1 and setup-2. In setup-1, all FPGA endpoints under the two T0s are active while in setup-2, 40 out of 64 FPGA endpoints are active.
We present results from both of these configurations as we observe small unexpected performance variations depending on which and how many switch ports are active in experiments near the peak network performance levels using all switch ports (setup-1). These variations appear to be related to internal switch microarchitectural details such as port-buffer affinity and vendor-specific scheduling policies. The setup-2 uses a subset of the switch ports and eliminates most of these vendor-specific and implementation-related behaviors. To get a better understanding of this behavior we performed a sweep where we capped the TX rate of our FPGA NICs and discovered that the slight degradation for setup-1 when using ARCANE appears to subside if the TX rate is capped at 95 Gbps.
@@ 4.4.2. Asymmetric Network Conditions
We evaluate the performance of ARCANE under asymmetric network conditions. We connect 16 endpoints through two T0 switches (8 endpoints each) with a total of 4 links to a pair of T1 switches. To demonstrate the adaptive load balancing capabilities of ARCANE, we change the link speed of one T0-T1 link from 400 Gbps to 200 Gbps, creating asymmetry in the network. Figure 9(b) shows the per-flow goodput as observed by the application while Figure 10(a) the FCT distribution. OPS sends packets across all paths (including those crossing the 200 Gbps link) with equal probability and is ultimately capped by the slower 200 Gbps path. The ECN marking on the 200 Gbps path causes the CC algorithm to throttle all flows and eventually match the capacity of that single slower link, thus leading to underutilization of the remaining 400 Gbps links (that are running at 50% utilization).
ARCANE can gracefully adapt in such a scenario as the cached entropies will reflect the network asymmetry and result in a path distribution that is skewed to tailor to the relative capacity of the available paths. In this example, ARCANE can reach high utilization with average per-flow goodput within 5% of the ideal fair-share target.
@@
Figure 9. ARCANE-FPGA impact on goodput.
@@
Figure 10. ARCANE-FPGA impact on FCT and packet drops
@@ 4.4.3. Network Failures
We also evaluate the performance of ARCANE in the presence of network failures. To demonstrate the robustness and resilience of ARCANE in the context of link failures, Figure 10(b) shows total packet drops (average across five runs each) observed in a large-scale 128 endpoint run (endpoints split across 2 T0s connected through 8 T1s) where we abruptly bring down a T0-T1 link during the experiment. While the network is trying to recover from the impact of this event (which in our environment can take in the order of 100s of milliseconds), OPS continues sending packets across all paths (including those affected by the link that went down). The freezing capability of ARCANE can quickly adapt to such events (within the order of an RTO) and avoid sending packets down the affected paths as the entropy cache is replenished from packets traversing healthy paths.
4.5. ARCANE Applicability
In this section, we briefly evaluate ARCANE, in simulations, under different scenarios by changing the ACK coalescing ratio, the EVS size and, the underlying CC algorithm. Generally, we believe that ARCANE can work well under many different circumstances, topologies and workloads.
4.5.1. ACK Coalescing
We have primarily evaluated ARCANE without ACK coalescing, as this configuration allows ARCANE to operate with the most up to date data. However, some transport protocols permit ACK coalescing, where the receiver sends an ACK packet only after receiving $n$ data packets from the sender. In theory, the coalesced ACK packet could return all previous non-ECN marked entropies in its header, but we focus on the worst case scenario for ARCANE where the ACK only carries the entropy of the packet that triggered it.
We run a real 8 MiB permutation workload simulating different ACK coalescing ratios. As we can see in the left part of Figure 11, ARCANE with 2:1, 4:1 and 8:1 coalescing ratios does significantly better than OPS, while with 16:1 it starts losing its advantage. However, it should be noted that in case of asymmetries or failures (right Figure 11), ARCANE does much better due to its various improvements over OPS even at 16:1 ratio. We confirm these results theoretically in Appendix D.1.
Finally, we note that the trade-off of sending more ACKs is worth the effort if the underlining hardware supports such rate and the impact on the network traffic is minimal (~1%) since ACKs are relatively small (64B) compared to the packet size usually used in modern interconnects (4KiB or 8KiB).
@@ 4.5.2. EVS Size
In Appendix B, we prove that OPS routing, done through EVs and ECMP hashing, requires a large EVS to work correctly and reduce collisions. Moreover, this requirement also grows with the number of output ports in a switch as demonstrated in Section 5.1. On the other hand, ARCANE, due to its adaptive nature, can drastically reduce the EVS size and still work well. We show this in the left plot of Figure 12 where we compare, in a real scenario, OPS and ARCANE when using 32, 256, and 64K EVs. ARCANE works equally well with 256 and 64K EVs and is only 8% slower with 32 EVs. On the other hand, OPS is 21% and 64% slower with 256 and 32 EVs when compared to 64K. This confirms that ARCANE could potentially work well even just with 1 byte for the EVS. We note that while OPS could be implemented without using EVS—such as through round-robin selection or by making random choices directly at the switch—these approaches introduce additional challenges (Section 5.2).
@@ 4.5.3. Different CC Algorithms
ARCANE has been designed to work with any CC algorithm as long as there is no over-reaction to out-of-order packets and ECN support. In this paper, we have, so far, evaluated ARCANE working alongside a tuned version of DCTCP. However, as we will see, ARCANE can work well even with other algorithms. Moreover, we envision also a version of ARCANE that could work just with delay if ECN is not supported but we do not go into details here. For example, in the right plot of Figure 12, we run a 8 MiB permutation workload without failure for DCTCP, EQDS and a proprietary CC algorithm. ARCANE can help all of these CC algorithms when compared to OPS.
@@ 5. Theoretical Verification
To support our experimental findings, we present a theoretical first principles analysis of OPS and demonstrate how it can lead to arbitrarily long queues. To address this issue, we develop a new recycled balls-into-bins model and prove its local convergence. This model serves as the theoretical intuition behind the ARCANE protocol (Section 3).
@@ 5.1. Recycled Balls-into-Bins Model
At maximum injection rate, OPS suffers from severe load-imbalance that eventually leads to queues of arbitrary length building up. We explain this behavior with an infinite batched balls-into-bins model (Berenbrink et al., 2018; Becchetti et al., 2019; Los and Sauerwald, 2023). In contrast, we show that recycling good paths such as in ARCANE leads to a convergent behavior and logarithmically-bounded queues.
@@
For $n\geq 16$, $\tau\geq 4\ln n$, $b\geq 2.4\ln n$, recycled balls-into-bins converges in $O(n\log n)$ steps. Every bin has $O(\log n)$ elements throughout, with probability $1-o(1)$.
@@
To visualize these behaviors, we model the balls-into-bins problem with both OPS and with the recycled balls improvement. In Figure 14, we set $n=5$ (for visualization purposes, but this holds true for more realistic $n$ values) and showcase the queues’ evolution for OPS and the recycled balls model. The results confirm that while OPS leads to unbounded queue growth, the recycled balls model converges, keeping all queues below the threshold ($\tau$) value, consistent with our previous network simulations (e.g., Figure 1 and Figure 3).
@@ 5.2. Limitations and Alternatives
While the provided model performs well in an idealized setting with maximal injection rates, in a real network scenario, CC algorithms would reduce the sending rate if a large queue were to build up. However, such activations of the CC would result in a slowdown of the sending rate and, hence, an increase of the completion time of a workload as we can see in Section 4.3.1. Moreover, we note that while real scenarios would have more subtle interactions not modelled here (like using distinct $KMin$ and $KMax$), the key results hold and are consistent with our evaluations in Section 4. Finally, we note that here we provide a proof only about the local convergence of ARCANE and that other models such as static path assignments and round-robin across the ports would also work well and potentially provide no queueing. However, such approaches are not viable in reality for a series of reasons: they cannot react well to partial or total failures, they would struggle with multi-tier topologies and they require detailed knowledge of the underlying workloads (Gangidi et al., 2024).
@@ 6. Related Work
The literature for load balancing designs is vast, however most of it focuses on TCP-based deployments where out-of-order packets are not desirable (Benson et al., 2011; Zhang et al., 2017; Al-Fares et al., 2010; Vanini et al., 2017; Dixit et al., 2013). In contrast, ARCANE aligns with a new transport paradigm (UltraEthernet (Consortium, 2024), Adaptive RoCE by NVIDIA (NVIDIA, 2024), Falcon (goo, 2023)) where out-of-order packets help to leverage the capacity of multi-path networks.
@@
Load balancers with per-packet granularity, such as OPS (Dixit et al., 2013) or DRILL (Ghorbani et al., 2017), help drastically reduce ECMP collisions, but are still oblivious to asymmetries (for OPS) or require switch support (for DRILL). As demonstrated in Section 5 OPS can suffer even in symmetric cases. MPRDMA (Lu et al., 2018) also uses ECN for load balancing like ARCANE, however it requires probing and ACK clocking and does not offer caching of entropies. Hermes (Zhang et al., 2017) combines both ECN and delay but it works best with TCP-like protocols and has many parameters with complex tuning.
@@
Proteus (Hu et al., 2023) focuses on optimizing the load balancing of lossless PFC (Priority Flow Control) networks while ARCANE focuses on lossy networks.
@@ 7. Conclusion
We presented ARCANE, a simple, lightweight, yet highly effective load-balancing mechanism designed to meet the constraints of next-generation datacenter networks tailored for AI workloads. As demonstrated in our extensive evaluations, conducted through both simulations and FPGA hardware, ARCANE’ adaptive entropy caching enhances end-to-end performance across multiple critical metrics, including average flow completion time, runtime, and packet loss. ARCANE outperforms ECMP and OPS by up to 6x and 1.25x in symmetric networks, and by up to 4.5x and 2x in asymmetric networks, outperforming OPS by as much as 100x during short-term transient link failures while reducing packet drops by over 70x. Moreover, ARCANE consistently beats or matches other state-of-the-art algorithms in all scenarios evaluated. We also showed how ARCANE can work well under various network configurations demonstrating its flexibility to adapt to different scenarios while remaining light-weight on NIC memory requiring only 25 bytes of state per-connection.
Appendix A Freezing Mode in ARCANE
Ideally, ARCANE should enter freezing mode only upon detecting a network failure. To achieve this, we employ two strategies: